OpenStudy (anonymous):

s(t)= 3t^2+2t-2 a. find the velocity for time t b.find the acceleration for time t c. compute the total distance I derived the velocity equation and the acceleration as 5t+2 and 5 but the solution checker online says its wrong....

OpenStudy (anonymous):

for c the interval is [0,1]

myininaya (myininaya):

v=s'=3(2)t+2=6t+2

OpenStudy (anonymous):

a. first derivative b. 2nd derivative c. s(1)

myininaya (myininaya):

your number in front of t is a little off

OpenStudy (anonymous):

wow.... this is what studying physics for 30 hours does to you....

OpenStudy (anonymous):

how did i not see that rofl

OpenStudy (anonymous):

you guys probably are facepalming right now. "how is this kid in calc 1??"

OpenStudy (anonymous):

Hey can you do part c I think Im making another alegbra error and I cant see it I'm just not sure what to do with v=0 @ t=-1/3

OpenStudy (anonymous):

s(1)-s(0)???????

OpenStudy (anonymous):

as a thought jesus man you need more question marks?

OpenStudy (anonymous):

Im pretty sure you can read more than just numbers as I said above I have been studying physics for 30 hours.

OpenStudy (anonymous):

maybe take a break!

OpenStudy (anonymous):

breaks are for sane people

OpenStudy (anonymous):

s(t)= 3t^2+2t-2 v = s'(t) = 6t^2 + 2 a = v'(t) = 12t